OpenStudy (anonymous):

How do you simplify (3 times the square root of 5)^2

OpenStudy (anonymous):

(3 sqrt 5)^2 = 3^2 • sqrt 5^2 = 9 • 5 = 45

OpenStudy (akashdeepdeb):

(3 rt 5)^2 = 3^2 . (rt5)^ 2 Get it! So answer = 9 * 5!!